OpenStudy (anonymous):

If the sin 60° is square root 3 over 2, then the cos ____° = _____.

OpenStudy (anonymous):

30° ; 1 over 2 60° ; square root 2 over 2 30° ; square root 3 over 2 60° ; 1

OpenStudy (anonymous):

\[\sin 60° = \sin (90^0-30^0) = \cos 30^0 = \frac {\sqrt3}{2}\] @hayleynicole1

OpenStudy (anonymous):

Thank you so much! i've been stuck on this question for awhile.

OpenStudy (anonymous):

@hayleynicole1 medal please